OpenStudy (iconforhire56):

How many solutions does the equation 5m − 5m − 12 = 14 − 2 have?

OpenStudy (freckles):

do you know what 5m-5m is?

OpenStudy (iconforhire56):

-5?

OpenStudy (freckles):

if you have 5 apples and someone takes away all your 5 apples then how many apples do you have?

OpenStudy (iconforhire56):

0

OpenStudy (freckles):

so 5 apples-5 apples is 0 so if you have 5 amount of m's and someone takes away your 5 amount of m's then how many m's do you have left?

OpenStudy (iconforhire56):

0

OpenStudy (freckles):

5m-5m=0m aka 0

OpenStudy (freckles):

so your equation is 0m-12=12 or -12=12 is -12=12 true?

OpenStudy (iconforhire56):

yeah so there is no solution?

OpenStudy (freckles):

right -12 is not the same 12 there is no value of m that will satisfy 5m-5m-12=14-2

OpenStudy (freckles):

no solution is correct since there is no m to satisfy that equation